OpenStudy (anonymous):

In special relativity, to derive the four-velocity from the four-position, one generally needs to do something of this manner:\[\vec V = \frac{d}{dt} \vec X = \left(\frac{d}{d \tau} \vec X \right) \frac{dt}{d\tau}\]Here, \(\tau\) is the proper time of the moving object, and \(t\) is our measurement of time. Now, in every derivation that I've seen, \(\frac{dt}{d\tau} = \gamma = \frac{1}{\sqrt{1-v^2/c^2}}\). However, time dilation takes the factor \(\gamma\), so wouldn't we expect \(\tau = \gamma t\), which would make \(\frac{dt}{d\tau} = \frac{1}{\gamma}\)? What am I mixing up?

OpenStudy (anonymous):

THE DERIVATIVE EQUATION OUTSIDE THE BRACKET. THERE I THINK U HAVE TO WRITE \[d \tau/dt and \not dt/d\] . but even i m not sure

OpenStudy (anonymous):

*\[d \tau/dt\]

OpenStudy (anonymous):

not

OpenStudy (anonymous):

\[dt/d\]

OpenStudy (anonymous):

sorry for posting again and again but thought the eq i posted the first time was not clear

OpenStudy (anonymous):

OHH good catch LMAO... since I wrote \(\frac{dt}{d\tau}\) instead of \(\frac{d\tau}{dt}\), that would explain why I was getting the reciprocal answer. The correct simplification would be\[\vec V = \frac{d}{dt} \vec X = \left(\frac{d}{d \tau} \vec X \right) \frac{d\tau }{dt} = \boxed{\left(\dfrac{d}{d \tau} \vec X \right) \gamma}.\]Thanks for catching that silly mistake!

OpenStudy (anonymous):

we all do something like that at some point of time glad to help.